Find: ( (iii) 2.71 × 5 (iv) 20.1 × 4 nveahuyghuyhdsbvhvuefwqtwyt8734yu85h8gueqwgy289yu38huj8ehyu

Find:

(iii) 2.71 × 5

(iv) 20.1 × 4

 

nveahuyghuyhdsbvhvuefwqtwyt8734yu85h8gueqwgy289yu38huj8ehyu

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

(iii) 2.71 × 5

Solution:- We have,

= (271/100) × 5

= (1355/100)
On dividing a decimal by 100, the decimal point is shifted to the left by two places. Then,

= 13.55

(iv) 20.1 × 4

Solution:- We have,

= (201/10) × 4

= (804/10)
On dividing a decimal by 10, the decimal point is shifted to the left by one place. Then,



Thanks
Vijay Tiwari
2 years ago
1) 13.5
2)80.4
 You can use calculator and see the answers are correct or not........................................